OpenStudy (caerus):

an a-c source delivers current to two loads A and B that are connected in parallel. If that A load current of 25 A leads the B load current of 42.2 A by 45 elec deg. a. calculate the resultant current and b. determine the angle bet. the resultant current and an assumed A current reference. ans= 62.6 and 28.5deg lag

OpenStudy (caerus):

@osprey

OpenStudy (irishboy123):

same here, i think the first answer is slightly offf but close, the phase lag is fine. phasor it !

OpenStudy (caerus):

can you solve it for me @IrishBoy123

OpenStudy (caerus):

\[\sqrt{(42.2\cos(45)+25)^2+(42.2\sin(45))^2}\]

OpenStudy (caerus):

=62.4,, how can i get the angle?

OpenStudy (caerus):

@IrishBoy123

OpenStudy (irishboy123):